OpenStudy (anonymous):

PI = 3.1415926 , i got a huge shock when i noticed this value , ever since ma 4th grade i knew pi=22/7 that gives pi= 3.142857 , but 22/7 is not actual pi, its an approximation , GROSS!!

pooja195 (pooja195):

ok?

OpenStudy (anonymous):

you didnt get me or did you lol

pooja195 (pooja195):

something about pi

OpenStudy (anonymous):

lol

OpenStudy (noelgreco):

22/7 is rational, and therefore a repeating decimal. Pi is irrational. I wish you'd been my student.

OpenStudy (anonymous):

Even what you have here, PI = 3.1415926, is still an approximation....

OpenStudy (anonymous):

@NoelGreco pi can never be expressed as rational number and @dpaInc can you please tell me from where does the value originate

OpenStudy (anonymous):

i don't understand ur question... but pi is basically the RATIO of a circle's circumference to its diameter: \(\large \pi = \frac{Circumference}{diameter} \) Wait.... ^^^ that's a RATIO of two numbers.... hmmm... maybe either circumference or diameter must be irrational also...

OpenStudy (anonymous):

i knew pi=22/7 , recently i studied an article n gotta know that 22/7 is an approximation n like that ther r other numbers now ma question is from where does the value of pi originate, k lets say i have a circle of dia =2 now ratio of dia and circumference would be = pi, but what gives value of pi = 3.1415? where does this value originate from

OpenStudy (anonymous):

sorry... i still don't understand ur question... that ratio just means if you take a piece of string the length of the diameter of ANY circle, that string will fit into the circumference slightly more than 3 times.... about 3.14 times.... and this is true for ANY circle.

OpenStudy (anonymous):

k lets make it simple for one more time, concept of pi ??? and how pi equals to 3.1415.... ?

OpenStudy (anonymous):

pi IS the constant RATIO of circumference to diameter... that ratio is the same no matter the circle u look at.

OpenStudy (anonymous):

m sorry still i am just wondering fromwhere does pi = that constant number, what is the calculation of that constant

OpenStudy (noelgreco):

...and how accurately did you measure the circumference and diameter? Tell me your precision, and I can be more precise. There is a proof for the irrationality of pi here: http://www.coolissues.com/mathematics/Pi/pi.htm

OpenStudy (anonymous):

fellas i understand all of you, i know you guys tryna say everything m just wondering from where does this constant originate , lets say planck's constant, it has been derived, lez say permittivity permeability every constant has been derived but where does pi come from

OpenStudy (anonymous):

what gives value of pi = 3.1415.....? where does pi originate from? i know it could be a waste to talk about it but this constant has been ma favorite since 4th grade n now it has drawn a huge attention

OpenStudy (anonymous):

where does c, speed of light, come from?

OpenStudy (noelgreco):

\[\frac{ \pi }{ 4 }=1-\frac{ 1 }{ 3 }+\frac{ 1 }{ 5 }-\frac{ 1 }{ 7 }....\] Keep adding terms for additional accuracy

OpenStudy (anonymous):

C has a very concrete solution in modern physics :) and @NoelGreco this is ma question where does this generate , idea of pi?

OpenStudy (anonymous):

and what is it? c is the constant speed of light.... pi is the constant ratio of circ/diameter

OpenStudy (anonymous):

k m tryna find where does pi come from but c has a whole derivation ;)

OpenStudy (anonymous):

good luck guy... :)

OpenStudy (anonymous):

:)

OpenStudy (accessdenied):

I imagine that pi originiates, in a sense, from curiosity of some of the earliest mathematicians/geometers. They would see that all these circular things look similar, and so maybe there's a common theme. The easiest things to measure are circumference and diameter, so that'd be a natural step. What they could do, then, is look for patterns with their results. Does any ratio appear by messing with our numbers? Earlier math worked with fractions and no calculators, though, so they had to do it the long way. They would eventually get some approximations going, 22/7 could have been one. I can't say that's exactly what happened, its a little bit of speculation. Of course, as math evolves, different approaches would come up like using Calculus / series. (: --- An interesting video by numberphile (a nice video series on cool numbers): https://www.youtube.com/watch?v=FpyrF_Ci2TQ Might be interesting for you as well.

OpenStudy (anonymous):

@AccessDenied thank you very much :)